### v1

# A Fully Adaptive Radau Method for the Efficient Solution of Stiff Ordinary Differential Equations at Low Tolerances

Abstract
Radau IIA methods, specifically the adaptive order radau method in Fortran due to Hairer, are known to be state-of-the-art for the high-accuracy solution of highly stiff ordinary differential equations (ODEs). However, the traditional implementation was specialized to a specific range of tolerance, in particular only supporting 5th, 9th, and 13th order versions of the tableau and only derived in double precision floating point, thus limiting the ability to be truly general purpose for highly accurate scenarios. To alleviate these constraints, we implement an adaptive-time adaptive-order Radau method which can derive the coefficients for the Radau IIA embedded tableau to any order on the fly to any precision. Additionally, our Julia-based implementation includes many modernizations to improve performance, including improvements to the order adaptation scheme and improved linear algebra integrations. In a head-to-head benchmark against the classic Fortran implementation, we demonstrate our implementation is approximately 2x across a range of stiff ODEs. We benchmark our algorithm against several well-reputed numerical integrators for stiff ODEs and find state-of-the-art performance on several test problems, with a 1.5-times speed-up over common numerical integrators for stiff ODEs when low error tolerance is required. The newly implemented method is distributed in open source software for free usage on stiff ODEs.
I Introduction
High-precision solving of ordinary differential equations (ODEs) is required in many disciplines, from controls and safety scenarios such as identifying potential asteroid collisions [1], to robustly resolving chemical reactions in astrochemistry which span more time scales than which 64-bit floating point numbers can account for [2]. There are many traditional scenarios where high precision is required as well, for example, in pharmacometrics it is a requirement during inference to solve the model to a much lower tolerance than what’s generally used in simulation (tolerances of around 1e-10) in order to improve the robustness of the fitting schemes [3]. This is due to the fact that calculating derivatives of ODE solutions incurs error from the forward pass and induced solver error from the extended equations for the forward/adjoint pass, making higher forward simulation accuracy a requirement in order to achieve accurate gradients and thus a stable optimization scheme [4] [5].
There is a wealth of high-order schemes for non-stiff ODEs, including the higher order Verner methods [6], dop853 [7], high order Gauss-Legendre schemes [8], and adaptive-order adaptive-time Adams-Bashforth-Moulton schemes like those developed in LSODE [9] and CVODE [10] [11]. Higher order schemes tend to be more efficient at computing solutions to higher accuracy due to the faster convergence with respect to shortened $dt$ [7]. In the SciML open source benchmarks, on non-stiff equations at high accuracy (required errors of roughly 1e-10 and below), schemes above order 5 such as Vern9 typically dominate in terms of performance.
However, there are many fundamental difficulties for achieving higher-precision solving via high-order methods for stiff equations. Importantly, it is known for many of the standard schemes that an order of 5 is the absolute maximum that can be achieved. For example, with implicit backwards differentiation formulae (BDF) integrators it is known that there is an order barrier where the 6th order method is too unstable for most use cases and all orders above are not zero stable [12], limiting all standard implementations such as lsoda [13], CVODE [10] [11], ode15s [14], to be an adaptive order scheme where order cannot go above 5 (but is generally bounded below 3 for highly stiff equations for stability region concerns). Meanwhile, singularly-implicit diagonal Runge-Kutta schemes (SDIRK) and common Rosenbrock-Wanner schemes with single $\gamma$ all have a maximum order of 5 [11], which is the reason the vast majority of other common stiff ODE solvers are below this order bound such as rodas [15], the FATODE suite [16], Rodas5P [17], and many more. Implicit extrapolation methods could use extra internal stages to overcome this order barrier and achieve arbitrary order, though traditional Fortran implementations such as Seulex and SODEX were never able to achieve the efficiency of the more commonly used ODE solver schemes and the recent Julia advances require using extra compute (exploiting multithreading) in order to achieve state-of-the-art performance [18].
Thus the only commonly used stiff ODE software for stiff equations which overcomes this order barrier are the RadauIIA schemes, which were popularized by Ernst Hairer’s implementation of radau in Fortran [19]. This implementation was a major achievement which uses the fact that RadauIIA schemes of arbitrary order can be derived and are A-B-L stable, giving a platform to build an arbitrary order scheme which is not stability bound for the higher order calculations. However, in order to reduce the computation and due to fundamental programming limitations of the time, the radau scheme pre-derived the schemes for only order 5, 9, and 13 and only for double precision (64-bit) floating point numbers. As such, the scheme tends to perform well for problems defined in double precision which require sufficient accuracy, doing well in the SciML benchmarks for problems which range between 1e-10 to the 64-bit floating point accuracy limit of around 1e-14. In addition, this older limitation does not effectively exploit many of the features of modern hardware, such as pervasive multithreading, single-input multiple data (SIMD) registers, and higher-precision / mixed-precision arithmetic.
For these reasons, we developed a modern implementation of the adaptive-order adaptive-time RadauIIA schemes in Julia which overcomes the previously mentioned issues and is able to fully exploit modern hardware. The software we have developed is available for use in Julia’s OrdinaryDiffEq.jl package, part of the DifferentialEquations.jl interface [24]. Much of the theory detailed in this article is adapted from Hairer’s original implementation [7] and [11]. However, our modernization includes a few crucial details, such as the ability to automatically derive order coefficients to arbitrary order to arbitrary user-chosen precision. As such, our adaptive-order scheme can automatically adapt to any order limit chosen by the user, and ultimately generate efficient schemes to any error tolerance provided by the user. Through benchmarks we demonstrate that this scheme achieves state-of-the-art efficiency when highly accurate solutions to stiff ODEs are required.
I-A Runge-Kutta Methods
Runge-Kutta methods are generalized methods to numerically approximating solutions to first order differential equations (ODEs) [20]. They are collocation methods that use intermediate approximations of the solution to extrapolate the overall solution.
Consider an initial value problem as follows:
$\frac{\mathrm{d}y}{\mathrm{d}t}=f(t,y)$ (1)
$y(t_{0})=y_{0}$ (2)
Such a problem can numerically be approximated at some future point in time $t$ using a Runge-Kutta method. The number of stages $s$ of a Runge-Kutta method denotes the number of nodes at which to take approximations of the solution. For an $s$ stage Runge-Kutta method with step size $dt$, the solution to $(1)$ at $t$ can be given as:
$y(t)=y_{0}+dt\cdot\Sigma_{i=1}^{s}{b_{i}k_{i}}$ (3)
Where $k_{p}$ is defined as
$k_{p}=f(t_{n}+c_{p}\cdot dt,y_{n}+\Sigma_{i=1}^{p-1}{a_{p,i}k_{i}})$ (4)
for $1\leq p\leq s$.
The constants $a_{i,j}$ for $1\leq i,j\leq s$, $b_{i}$ for $1\leq i\leq s$, and $c_{i}$ for $1\leq i\leq s$ are typically represented by a Butcher tableau, as shown below:
$\begin{array}[]{c|ccccc}c_{1}&a_{11}&a_{12}&\cdots&a_{1,s}\\ c_{2}&a_{21}&a_{22}&\cdots&a_{2,s}\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 1&a_{s,1}&a_{s,2}&\cdots&a_{s,s}\\ \hline\cr&b_{1}&b_{2}&\cdots&b_{s}\end{array}$
$a$ is defined as the internal coefficients, $b$ is defined as the weights, and $c$ is defined as the nodes [7]. Throughout this paper, we take $c_{s}=1$ as is defined for a RadauIIA method.
We define the order of a Runge-Kutta method to be the maximum number $n$ such that the Taylor series expansion of $y(t_{0}+dt)$ and $y_{1}$ coincide for all terms from $1$ to $n$ [7].
The constants $a_{i,j}$, $b_{i}$, and $c_{i}$ typically satisfy some order conditions that make them more usable. For example, the first order conditions for a Runge-Kutta Method is
$\Sigma_{i=1}^{s}{b_{i}}=1$ (5)
While the second order condition is:
$\Sigma_{i=1}^{s}{b_{i}c_{i}}=\frac{1}{2}$ (6)
Such order conditions extend to third order conditions and further, and are important for testing whether a given method truly matches its assigned order. Higher order methods should approximate the solution faster as the timesteps shrink, so these order conditions are important in checking whether the method is of the appropriate order. The derivations of the above definitions can be found in [7].
Several further ”simplifying assumptions” are commonly used, for example:
$\Sigma_{j=1}^{i-1}{a_{ij}}=c_{i}$ (7)
but this condition is neither necessary nor sufficient for consistency. Many similar conditions exist; they merely serve to simplify the derivation of coefficients [7].
Another important property of Runge-Kutta methods to consider is their stability. The stability of a numerical method describes the tendency of local errors to either be amplified or diminished across time-steps. In other words, the stability properties of a Runge-Kutta method determines whether or not the method will converge globally as $dt$ goes to 0 [11].
The stability function of a Runge-Kutta method is found when the method is applied to an ODE of the form $y^{\prime}=\lambda y$. When Runge-Kutta methods use selected $dt$ that fall in the stability region, they prove to be highly accurate at approximating the true solution [11].
I-B Stiff ODEs
A subset of ODEs, known as stiff ODEs, pose particular challenges for most numerical integrators. A stiff ODE is one where there exists a dissonance between a slow and fast phase in the ODE, where the slow phase needs to be resolved while avoiding divergence in the fast phase.
As an example of a stiff ODE, consider the well known chemical reaction known as the Robertson equation [11]:
[…]
When translating this system into a system of linear ODEs, there will be a clear difference in the rates of equations 8 and 9, which will lead most Runge-Kutta methods struggle to handle the contrast between these two phases. An efficient numerical integrator will take as large time-steps as is possible to approach the solution quickly, but doing so in this case would cause the method to diverge from the true solution because of the rapid changes in equation 9. It is here where we see that stability of a numerical integrator is crucial, as having a wider range of options for the step-size allows the method to be more efficient on these systems [11]. As we will see in the next section, Radau methods posses stability properties necessary to overcome these challenges.
I-C Radau Methods
Radau methods are a class of implicit Runge-Kutta methods that are defined by quadrature formulas. An explicit Runge-Kutta method is one where the derivation of the current state is only done using previous states [11]. The simplest example is the explicit Euler method:
$y(t_{0}+h)=y_{0}+dt\cdot f(t_{0}+dt,y_{0})$ (11)
However, implicit Runge-Kutta methods use an implicit definition of the current state in addition to past states in deriving the current state, as seen below, in the implicit Euler method:
$y_{1}=y_{0}+dt\cdot f(t_{1},y_{1})$ (12)
When discussing the benefits of using Radau methods, it’s crucial to consider their stability, which is one of their prize-properties. For solving stiff ODEs, strong stability properties are crucial to ensure that the numerical method can handle the perturbations [11].
Radau methods have those stability properties. They are A-stable methods, meaning that when they are applied to an ODE of the form $\frac{dy}{dx}=qx$, where $q$ has a negative real part, the solutions will trend to 0 for all positive $dt$ [11]. Graphically, this property means that our method is stable on the entire left half of the complex plane [11]. That large stability region allows Radau methods to converge across a large range of time-steps, an important property while trying to solve stiff ODEs.
Furthermore a more crucial property for solving stiff ODEs is L-stability, which a method posses when it is both A-stable and has the stability function $g(dt)$ go to 0 as $dt\rightarrow\infty$ [11]. This property makes the method robust to the oscillations characteristic of stiff ODEs, and Radau methods also have this property. As a result, they are among the most efficient numerical integration methods for solving stiff ODEs.
Consider the below comparison between one of the most popular explicit methods, Dopri15, and RadauIIA5, a 5th order Radau method. The difference in performance is glaring. Radau, with its bigger stability region and more optimal properties, quickly converges on the solution. However, Dopri, struggling to resolve the perturbations in the equation, fluctuates around the true solution.
Figure 1: Comparison between different Runge-Kutta methods on a stiff ODE (source: [11])
In this paper, we describe our implementation of a Radau method that builds on the existing art, increasing the speed at which stiff ODEs can be solved.
II The State of Radau and our Developments
The most popular Radau software was developed by Ernst Hairer and published in his landmark paper “Stiff differential equations solved by Radau methods”. There, he laid out his implementation of a Radau method that autonomously switches between methods of 5th, 9th, and 13th order [19]. By changing the order of the method, his algorithm can more appropriately adapt to challenges in solving the stiff ODE. If a certain part of solving an ODE requires smaller time-steps, the method can switch to a more efficient higher-order method. Likewise, if a method can take larger time-steps in another region of the solve, a lower order method helps to streamline the computation. So, by having an adaptive order method, the method can adjust with the problem. This allows the method to optimize both solve time and accuracy.
However, there were some limitations to the existing algorithm:
  1. 1.
    The implementation required hard-coded tableau coefficients.
  2. 2.
    The implementation was limited to methods of 5th, 9th, and 13th order, constraining the flexibility of the method.
A later paper from Martín-Vaquero adds a 17th order method to Hairer’s initial implementation, and sees performance boosts at low tolerances [21]. This indicates that adding further adaptive capabilities to the method will continue to improve on performance, motivating our work.
In our new adaptive Radau method, we address both of the limitations of Hairer’s initial work, thus enhancing the performance of the method.
III Implementation of the New Radau Method
III-A Programming Language Choice
We chose to use Julia to implement our new adaptive Radau method, a contrast to the existing implementation in Fortran. Julia is a high-level programming language which has a simplified interface (garbage collection, built-in array utilities, etc.) like Python or R but has an optimizing compiler and type inference as part of its runtime which allows for its code to be as fast as languages like C or Fortran [22]. This is important for applications such as ODE solving because one of the major bottlenecks to ODE solver performance is the performance of a user’s function definition, which when using a high-level programming language like Python will be unoptimized [23]. Additionally, the DifferentialEquations.jl has many advantages over traditional ODE solver libraries, such as support for arbitrary precision arithmetic, automatic differentiation, and automatic sparsity detection which allows for simple user code to achieve maximum performance and precision [24] [25] [26]. Our new adaptive Radau methods make use of all of this infrastructure in tandem, using the arbitrary precision arithmetic to support more accurate convergence with higher orders, automatic differentiation for faster and more accurate Jacobians, and automated sparsity detection as a feature for improving nonlinear solver performance when the underlying $f^{\prime}$ is sparse.
III-B Tableau Elements
In building the new adaptive Radau method, we developed an algorithm that autonomously calculates the tableau for a method of any given order. The derivation comes from the order conditions and simplifying assumptions, as shown below [21]:
$c$ is defined as the roots of the Radau polynomial:
$\frac{\mathrm{d}^{s-1}}{\mathrm{d}x^{s-1}}x^{s-1}(x-1)^{s}$ (13)
${a_{i,j}}$ is then defined as:
$\Sigma_{j=1}^{s}{a_{i,j}c_{j}^{q-1}}=\frac{c_{i}^{q}}{q},i=1,2,\cdots,s,q=1,2,\cdots,s$ (14)
$b$ is simply defined as:
$b_{i}=a_{s,i}$ (15)
Mathematically, this derivation would be sufficient for our Radau method. However, in implementing the method in code, we transform our computations into a basis where $a^{-1}$ has a defined form, because $a^{-1}$ appears while solving a costly linear system. Knowing the form of $a^{-1}$ helps streamline computation.
III-C Transformation Matrix
To understand the transformation matrix $T$, it is first important to understand the form of $a^{-1}$. Because we only operate on methods with an odd number of stages, $a^{-1}$ always has one real eigenvalue and several complex-conjugate pairs of eigenvalues. For example, when $s=3$, below is $a^{-1}$ and its eigenvectors:
$a^{-1}=$ $\begin{pmatrix}0.197&-0.066&0.233\\ 0.394&0.292&-0.042\\ 0.376&0.512&0.111\end{pmatrix}$
$\lambda_{1}=3.6378342527$
$\lambda_{2}=2.681082873627+3.0504301992474i$
$\lambda_{3}=2.681082873627-3.0504301992474i$
The goal now is to find $T$ such that we know the structure of $a^{-1}$ (diagonal, block-diagonal, etc). Naively, we may just take $T$ as the eigenvectors of $a^{-1}$, producing the diagonalization of $a^{-1}$, but this would yield a complex transformation matrix and a complex form of $a^{-1}$, which is suboptimal. Instead, for the below eigenvectors of $a^{-1}$:
$\begin{pmatrix}0.091&-0.128+0.027i&-0.128-0.027i\\ 0.242&0.186+0.348i&0.186-0.348i\\ 0.966&0.909&0.909\end{pmatrix}$
By letting $r$ be the real eigenvector, the complex eigenvectors be represented as $u+vi$ and $u-vi$, where $u$ is a vector holding the real components and $v$ is a vector holding the imaginary components of the complex eigenvectors, we can take a basis ${r,u,v}$, and construct our transformation matrix $T$. For simplicities sake, we also scale $u$ and $v$ to have either a $1$ or a 00 as their last element, streamlining the computation. Applying this transformation to $a^{-1}$ gives a block diagonal form for $a^{-1}$, with the complex eigenvalues represented as rotation matrices, as shown below:
$T=$ $\begin{pmatrix}0.09&-0.14&-0.03\\ 0.24&0.20&0.38\\ 0.96&1.0&0\\ \end{pmatrix}$
$T^{-1}a^{-1}=$ $\begin{pmatrix}3.64&0&0\\ 0&2.68&-3.05\\ 0&3.05&2.68\end{pmatrix}$
Thus, applying our transformation matrix $T$ allows us to have a sparser structure of $a^{-1}$, which can be exploited to split the LU-factorizations and thus greatly decrease the computational cost of the linear algebra.
III-D Embedded Set Calculation
A second element that our script derives is an adaptivity set. This adaptivity set is used to vary the step-size of our method. In regions where there is a low error-estimate for our method, we can take larger time- steps so as to optimize our solve time, while if there is a higher error-estimate, we take smaller time-steps to prevent divergence of our method.
The calculation of our time-steps depends on an error-estimate that is calculated based on the difference between our method and a lower order ”embedded method”. This method shares the tableau coefficients $a$ and $c$ with the original method, while having a modified $b$ set that allows the embedded method to have a lower order. The calculation of the $b$ set is done using the order conditions.
By taking the difference between the value calculated by the embedded method and our Radau method, we have an estimate of the error of the Radau method. Then, our predictive controller uses that error estimate to determine the optimal size for the next time-step, as will be described in the next section.
Our embedded set is calculated by considering our method as a collocation method [27], which then gives us the following linear system of equations that is easily solved for $b$:
$\Sigma_{j=1}^{s}{b_{i}c_{i}^{k-1}}=\frac{1}{k},k=1,2,\cdots,s$ (16)
III-E Tableau Caching
In our implementation, we make use of a dictionary to store the tableaus. This helps to alleviate some computation time and avoids repeated re-derivation of commonly used tableaus. However, it remains true that our method can calculate the coefficients for a method of any order on-the-fly. Caching the common tableaus merely helps to speed up the process for those most frequently used orders, like 5th, 9th, and 13th. If the method chooses to go beyond those orders, it will derive the tableau autonomously, cache the tableau, and continue the numerical integration process.
III-F Step-Size Adaptivity
The algorithm makes use of the standard predictive controller and step-size selection algorithm.
The general idea is to use our error estimate to determine whether our current step-size is too big or small, and to scale the size accordingly. A high error estimate indicates that we should take smaller time-steps, and vice-versa for low error estimates.
The controller changes $dt$ based on the error. The number of iterations needed to converge is also factored into the algorithm, as this is indicative of the convergence rate. Mathematically, we calculate parameters $f$ and $expo$ to determine the scaling factor, $q$.
$f=\frac{(1+2*max\_iter)*\gamma}{num\_iters+2*max\_iter}$ (17)
$expo=\frac{1}{(adaptive\_order+1)}\$ (18)
$q=\frac{error^{expo}}{f}$ (19)
Where $\gamma$, $max\_iter$, and $num\_iters$ are parameters for a given iteration of our solve. Then, we find our scale factor $q$, and multiply $dt$ by this value to determine what the next time-step should be [28].
III-G Order Adaptivity
In addition to step size adaptivity, our method also builds on the order adaptivity in existing methods by implementing a novel order selection algorithm. Where the existing implementation uses the value of the divergence to determine whether to raise or lower the order [19], our algorithm chooses a different approach. However, the general scheme is similar.
The time-steps of a Radau method uses several Newton iterations while converging on the solution to that timestep. So, we use a metric based on the number of iterations that it takes the method to converge. If the method is taking several iterations to converge, this implies that a lower order method will be more efficient.
[…]
  1. 1.
    If $\kappa<2.75$, then the method is converging sufficiently fast. Raise the order of the method by 4.
  2. 2.
    Else if $\kappa>8$, then the method is converging too slowly. Decrease the order of the method by 4.
  3. 3.
    Otherwise, continue at the same order.
Doing so allows us to combine step-size and order adaptivity. When we are taking small timesteps and need more iterations to converge, the algorithm chooses a higher order method, and vice versa. Thus, we optimize the solving of each section of the problem, thus optimizing the solve as a whole.
IV Results
To analyze the results of the method, we use work-precision diagrams, which show the time it takes a given solver to reach a certain degree of accuracy. Lower times and higher accuracies on the diagram indicate better performance. Accuracy is compared against solves of a given problem at tolerances of $10^{-14}$.
IV-A Competing algorithms
The algorithms against which we benchmark our adaptive Radau method are described below:
  1. 1.
    Hairer’s Radau: The existing implementation of Radau methods in Fortran.
  2. 2.
    BDF Methods: Another set of very common numerical integrators for stiff ODEs. They are multi-step methods that use interpolating polynomials to derive future time-steps. They initially struggled with stability issues, but later improvements have created very powerful methods that perform well on stiff ODEs [29].
  3. 3.
    Rosenbrock methods: A set of numerical integrators that perform only one iteration of Newton’s method [29]. Solving the linear system can be expensive in these methods, but when applied correctly, they are among the best numerical integrators for stiff ODEs [12][29].
BDF and Rosebrock methods cannot achive the same order of accuracy that Radau methods can. Therefore, while BDF and Rosenbrock methods tend to dominate solving stiff ODEs at high tolerances, Radau methods typically dominate the solution to low tolerances.
IV-B Work-Precision Diagrams
We test our adaptive Radau method against several methods belonging to the classes listed above and plot work-precision diagrams. The test problems, from Julia’s SciML benchmarks, are listed below:
  • Oregonator, 3 ODEs [30] [31].
  • Robertson, 3 ODEs [32].
  • Hires, 8 ODEs [33].
  • Pollution, 20 ODEs [34].
In each of the diagrams, the algorithm we implemented is labeled ”AdaptiveRadau”. We run the tests with the possible orders for the method as 5, 9, 13, 17, 21, or 25.
Figure 2: Comparison between different Runge-Kutta methods on the Oregonator problem
Figure 3: Comparison between different Runge-Kutta methods on the Robertson problem
Figure 4: Comparison between different Runge-Kutta methods on the Hires problem
Figure 5: Comparison between different Runge-Kutta methods on the Pollution problem
The comparison between only our AdaptiveRadau algorithm and the existing implementation of radau in Fortran is shown in the appendix, where results are consistently better by an average of 2-times across all test problems.
Furthermore, across all algorithms as a whole, we find that the AdaptiveRadau algorithm we implement achieves state-of-the-art performance on Hires, Oregonator, and Robertson problems. Even when the algorithm does not perform as well, like on the Pollution problem, we still far outperform the existing implementation of Radau.
We also consider benchmarks of our algorithm beyond the Float64 domain. In the following, we compare our AdaptiveRadau algorithm against one of the best stiff numerical integrators in Julia, Rodas5P, a newer high-order Rosenbrock method [17]. Both solutions are compared with a reference solution computed to tolerances of $10^{-24}$. The problem and solutions are computed with BigFloat precision.
Figure 6: Comparison between AdaptiveRadau and Rodas5P on the Oregonator problem at BigFloat precision
Figure 7: Comparison between AdaptiveRadau and Rodas5P on the Hires problem at BigFloat precision
V Discussion
Radau methods have dominated the solution of stiff ODEs at low tolerances since their development because they can achieve much higher orders [29]. In our work, we see that our algorithm continues to dominate that area. Our methods often are the fastest method to converge on the solution at low error, as seen in the work-precision diagrams above.
Radau methods had not been changed in decades, since Hairer’s implementation appeared to be the gold standard for high-precision stiff ODE solving. However, we now see that the standard has been raised by our algorithm. At high tolerances, we don’t outperform the best algorithms for solving stiff ODEs, but neither does Radau. Our implementation does compete with the Rosenbrock methods at these higher tolerances but is often outperformed. However, at low tolerances, the diagrams show a different story. Our algorithm is among the best, if not the best, at solving these stiff ODEs with high precision.
Furthermore, we show that the mixed-precision capabilities of our algorithm enhance the state-of-the-art. The comparison between Rodas and Radau on the BigFloat problem setups highlights the value of a method that can change orders. The relatively flat curve shown in our algorithm is starkly contrasted with the steeply sloping Rodas5P curve. By adjusting the order to the time steps taken, we optimize our solve. The diagrams show that we are extremely efficient at low tolerances, building on the success we had in the prior diagrams.
Acknowledgment
S.E. thanks the MIT PRIMES Computer Science program for connecting him with his mentor and giving him the chance to work on this research, and Dr. Chris Rackauckas and Oscar Smith for their continued support throughout the process.